car a has a headstart of 740 m ahead of car b car a is travelling at a constant 50km/h while car b starts from rest and accelerates at a constant 1.22m/s how long will it take car b to catch car a and how far will car b have traveled


1
Expert's answer
2021-10-28T08:51:37-0400

Distance of car A relative to car B:


"d=d_0+vt."

Distance of car B:


"d=at^2\/2."

Equate:

"d_0+vt=at^2/2,\\\space\\
t^2\frac a2-tv-d_0=0,\\\space\\
t=48\text s."

The distance car B traveled:


"d=at^2/2=1405\text m."